Solve the following problem:


A government office has six telephone lines. For the past months, the probability distribution of the random variable Y which represents the number busy line per day in shown in the records below



Y P(Y)


0 0.052


1 0.054


2 0.232


3 0.240


4 0.174


5 0.105


6 0.043



1. What is the probability that exactly four telephone


lines are busy in a day?


2.What is the probability that, at least ,four telephone lines busy in a day?


3.What is the probability that at least two but at most


four telephone lines busy in a day?


4.What is the probability that at least one telephone


lines are busy in a day?


5. What is the expected number of busy telephone lines in a day? Explain the results


6. What is the standard deviation of the number of busy telephone lines in a day? Explain the results.

1
Expert's answer
2022-02-22T03:31:25-0500

1) it is given in the condition. "P(Y=4)=0.174"

2) "P(Y\u22654)=P(Y=4)+P(Y=5)+P(Y=6)=0.174+0.105+0.043=0.322"

3)"P(2\u2264Y\u22644)=P(Y=2)+P(Y=3)+P(Y=4)=0.232+0.240+0.174=0.646"

4)"P(Y\u2265 1)=1-P(Y<1)=1-P(Y=0)=1-0.052=0.948"

5) "E(Y)=0*0.052+1*0.054+2*0.232+3*0.240+4*0.174+5*0.105+6*0.043=2.717"

This result means that in significant amount of time(month, for example), the mean number of busy lines per day will be around 2.717

6) "\sigma(Y)=\sqrt{E(Y^2)-E^2(Y)}=\sqrt{0^2*0.052+1^2*0.054+2^2*0.232+3^2*0.240+4^2*0.174+5^2*0.105+6^2*0.043-2.717^2}=1.65"

This result means that in significant amount of time(month, for example),we can expect that on around 68% of days the number of busy lines will deviate from the mean for no more than 1.65